40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4852     Entry time: Wed Jun 22 01:59:43 2011
Author: Sonali 
Type: Update 
Category: Green Locking 
Subject: fibre-coupling of the IR beam  

 What I did today.

1. Collimation of a beam.

  • I then practiced collimation of a 700 nm laser (output) beam after being coupled through a fibre.
  • I put together the set-up as shown in the attached picture where I used ....... to couple 650nm light into the PM.... fiber.
  • I kept shifting the focus of the output beam to an appreciable distance till it was approximately collimated.

2. Coupling of the IR light at the ETMY table to a fibre.

  • The fibre coupler was put in place to couple light into the fiber.
  • I put in the mirrors as planned to direct the IR beam exiting the doubling crystal towards the fiber coupler (input).
  • The mirrors were aligned such that the beam falls on the input lens of the coupler.
  • The far-end of the fibre originally would have gone to to PSL table but it has been put on this table to study the power of the IR beam transmitted through this set-up.  The output end of the fiber has been connected to another fiber optic coupler to collimate the exiting beam.
  • The picture of the current status attached.
Attachment 1: ETMY_june_21.jpg  3.596 MB  | Hide | Hide all
ETMY_june_21.jpg
Attachment 2: collimation_700nm_21_june.JPG  3.644 MB  | Hide | Hide all
collimation_700nm_21_june.JPG
ELOG V3.1.3-